a triangular garden has a perimeter of 53 yards. One side is 24 yards. The second side is 7 yards less than the third side. What are the lengths of the second and third side?
Answer by stanbon(75887)

a triangular garden has a perimeter of 53 yards. One side is 24 yards. The second side is 7 yards less than the third side. What are the lengths of the second and third side?
--------------
1st side: 24 yds
2nd side: x - 7 yds
3rd side: x
-----
Equation:
x + x-7 + 24 = 53
----
2x + 17 = 53
---
2x = 36
=======================
x = 18 yds (3rd side)
x-7 = 11 yds (2nd side)
=============================